OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]?

OpenStudy (cwrw238):

can you wok out what g(5) is?

OpenStudy (austinl):

You put 5 in for x on g(x). Then you put that value in for x in f(x).

OpenStudy (anonymous):

g(5)=24 f(g(5)=3*(24)-2 =72-2 =70

OpenStudy (anonymous):

thanks you all , I was really out of my mind with this